OpenStudy (anonymous):

What is the simplified form of 16q^5r^3 / 2q^4r^5 ?

jimthompson5910 (jim_thompson5910):

16/2 = ???

OpenStudy (anonymous):

8 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

q^5 divided by q^4 = ???

OpenStudy (anonymous):

Just q @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good, and r^3 divided by r^5 = ???

jimthompson5910 (jim_thompson5910):

Hint: If you're stuck, then think of  as